# Colorful Pinball: Density-Weighted Quantile Regression for Conditional Guarantee of Conformal Prediction

### v2

@@ A.5 Detailed results
Table 4: Comparison of MSCE. Mean Squared Coverage Error (MSCE, $\downarrow$) on 8 high-dimensional benchmarks.
@@
Split 0.0030 ± 0.0007 0.0127 ± 0.0029 0.0033 ± 0.0007 0.0349 ± 0.0120 0.0035 ± 0.0009 0.0073 ± 0.0008 0.0045 ± 0.0019 0.0123 ± 0.0017
PLCP (G=20) 0.0017 ± 0.0008 0.0022 ± 0.0032 0.0005 ± 0.0002 0.0248 ± 0.0142 0.0029 ± 0.0017 0.0015 ± 0.0011 0.0018 ± 0.0019 0.0024 ± 0.0022
PLCP (G=50) 0.0018 ± 0.0007 0.0022 ± 0.0027 0.0008 ± 0.0008 0.0172 ± 0.0100 0.0028 ± 0.0014 0.0014 ± 0.0007 0.0034 ± 0.0026 0.0016 ± 0.0012
Gaussian-Scoring 0.0010 ± 0.0005 0.0006 ± 0.0004 0.0005 ± 0.0002 0.0169 ± 0.0136 0.0021 ± 0.0028 0.0031 ± 0.0013 0.0020 ± 0.0019 0.0071 ± 0.0018
CQR 0.0015 ± 0.0010 0.0026 ± 0.0010 0.0006 ± 0.0004 0.0207 ± 0.0079 0.0013 ± 0.0005 0.0011 ± 0.0008 0.0013 ± 0.0013 0.0068 ± 0.0009
CQR-ALD 0.0012 ± 0.0010 0.0013 ± 0.0005 0.0006 ± 0.0003 0.0170 ± 0.0112 0.0010 ± 0.0005 0.0009 ± 0.0006 0.0010 ± 0.0005 0.0046 ± 0.0013
RCP 0.0016 ± 0.0005 0.0033 ± 0.0008 0.0006 ± 0.0003 0.0048 ± 0.0034 0.0013 ± 0.0003 0.0017 ± 0.0005 0.0009 ± 0.0005 0.0039 ± 0.0007
RCP-ALD 0.0010 ± 0.0004 0.0019 ± 0.0008 0.0004 ± 0.0002 0.0089 ± 0.0066 0.0011 ± 0.0004 0.0011 ± 0.0006 0.0007 ± 0.0003 0.0031 ± 0.0006
CPCP 0.0011 ± 0.0006 0.0006 ± 0.0003 0.0005 ± 0.0002 0.0020 ± 0.0011 0.0003 ± 0.0002 0.0008 ± 0.0004 0.0010 ± 0.0006 0.0015 ± 0.0005
CPCP (Mix) 0.0010 ± 0.0006 0.0005 ± 0.0003 0.0004 ± 0.0002 0.0021 ± 0.0013 0.0004 ± 0.0002 0.0008 ± 0.0003 0.0006 ± 0.0004 0.0011 ± 0.0003
CPCP (Clip) 0.0009 ± 0.0003 0.0004 ± 0.0002 0.0004 ± 0.0002 0.0019 ± 0.0008 0.0003 ± 0.0001 0.0008 ± 0.0003 0.0005 ± 0.0003 0.0011 ± 0.0004
CPCP (Clip+Mix) 0.0008 ± 0.0003 0.0004 ± 0.0002 0.0003 ± 0.0001 0.0017 ± 0.0008 0.0003 ± 0.0001 0.0007 ± 0.0004 0.0003 ± 0.0001 0.0010 ± 0.0003
Table 5: Comparison of WSC. Worst-Slice Coverage (WSC, $\uparrow$) on 8 high-dimensional benchmarks.
@@
Split 0.8083 ± 0.0265 0.6492 ± 0.0180 0.8134 ± 0.0183 0.5703 ± 0.0574 0.7470 ± 0.0097 0.7824 ± 0.0237 0.7239 ± 0.0194 0.7462 ± 0.0165
PLCP (G=20) 0.8593 ± 0.0286 0.8318 ± 0.0517 0.8775 ± 0.0176 0.6428 ± 0.0800 0.7709 ± 0.0494 0.8624 ± 0.0285 0.8150 ± 0.0472 0.8522 ± 0.0236
PLCP (G=50) 0.8534 ± 0.0242 0.8286 ± 0.0585 0.8711 ± 0.0218 0.6821 ± 0.0925 0.7803 ± 0.0441 0.8661 ± 0.0218 0.7936 ± 0.0499 0.8630 ± 0.0198
Gaussian-Scoring 0.8605 ± 0.0319 0.8607 ± 0.0140 0.8727 ± 0.0180 0.6723 ± 0.0870 0.8251 ± 0.0530 0.8314 ± 0.0239 0.8287 ± 0.0356 0.8017 ± 0.0224
CQR 0.8653 ± 0.0296 0.8162 ± 0.0166 0.8752 ± 0.0170 0.6440 ± 0.0560 0.8284 ± 0.0268 0.8682 ± 0.0224 0.8313 ± 0.0204 0.8050 ± 0.0162
CQR-ALD 0.8699 ± 0.0265 0.8457 ± 0.0148 0.8750 ± 0.0177 0.6724 ± 0.0708 0.8462 ± 0.0212 0.8753 ± 0.0196 0.8378 ± 0.0250 0.8121 ± 0.0195
RCP 0.8662 ± 0.0285 0.7952 ± 0.0155 0.8780 ± 0.0198 0.7963 ± 0.0546 0.8291 ± 0.0155 0.8572 ± 0.0223 0.8272 ± 0.0168 0.8318 ± 0.0117
RCP-ALD 0.8722 ± 0.0210 0.8310 ± 0.0173 0.8860 ± 0.0125 0.7459 ± 0.0827 0.8360 ± 0.0164 0.8709 ± 0.0217 0.8570 ± 0.0142 0.8428 ± 0.0139
CPCP 0.8757 ± 0.0189 0.8685 ± 0.0173 0.8824 ± 0.0180 0.8323 ± 0.0374 0.8886 ± 0.0145 0.8618 ± 0.0205 0.8599 ± 0.0135 0.8518 ± 0.0192
CPCP (Mix) 0.8840 ± 0.0235 0.8714 ± 0.0186 0.8842 ± 0.0161 0.8496 ± 0.0342 0.8800 ± 0.0125 0.8826 ± 0.0188 0.8689 ± 0.0139 0.8644 ± 0.0190
CPCP (Clip) 0.8887 ± 0.0194 0.8745 ± 0.0100 0.8834 ± 0.0147 0.8359 ± 0.0338 0.8842 ± 0.0175 0.8736 ± 0.0142 0.8690 ± 0.0133 0.8662 ± 0.0197
CPCP (Clip+Mix) 0.8932 ± 0.0142 0.8772 ± 0.0136 0.8857 ± 0.0126 0.8346 ± 0.0345 0.8833 ± 0.0138 0.8790 ± 0.0196 0.8805 ± 0.0124 0.8692 ± 0.0143
Table 6: Comparison of Size. Average Log Volume (Size, $\downarrow$) on 8 high-dimensional benchmarks.
@@
Split 0.7377 ± 0.0452 0.3924 ± 0.0101 0.1686 ± 0.0216 -1.1759 ± 0.5555 -1.9751 ± 0.0365 0.9245 ± 0.0292 -1.5225 ± 0.1065 1.2461 ± 0.0246
PLCP (G=20) 0.7303 ± 0.0618 0.3623 ± 0.0149 0.1286 ± 0.0185 -1.6360 ± 0.5384 -2.0241 ± 0.1109 0.8500 ± 0.0285 -1.6540 ± 0.1366 0.8960 ± 0.1054
PLCP (G=50) 0.7129 ± 0.0489 0.3592 ± 0.0172 0.1287 ± 0.0195 -1.7568 ± 0.4563 -2.0398 ± 0.0888 0.8417 ± 0.0283 -1.5798 ± 0.1398 0.8506 ± 0.0947
Gaussian-Scoring -0.7245 ± 0.0568 -1.6372 ± 0.0255 0.0378 ± 0.0176 0.4392 ± 0.7495 0.4323 ± 1.2047 -0.6169 ± 0.0347 -1.6076 ± 0.3728 1.2466 ± 0.2128
CQR 0.6460 ± 0.0595 0.2813 ± 0.0055 -0.0246 ± 0.0162 -1.0766 ± 0.2553 -2.5073 ± 0.0509 0.7827 ± 0.0181 -2.7842 ± 0.0691 0.6652 ± 0.0213
CQR-ALD 0.6657 ± 0.0616 0.2920 ± 0.0121 -0.0216 ± 0.0187 -0.8500 ± 0.3860 -2.6602 ± 0.0316 0.8497 ± 0.0325 -2.7822 ± 0.1037 0.9686 ± 0.0422
RCP 0.7531 ± 0.0699 0.3222 ± 0.0072 0.1240 ± 0.0288 -2.1207 ± 0.3476 -2.2085 ± 0.0510 0.8201 ± 0.0218 -1.8827 ± 0.0655 0.6992 ± 0.0367
RCP-ALD 0.7170 ± 0.0440 0.3321 ± 0.0083 0.1155 ± 0.0168 -1.9556 ± 0.4510 -2.2267 ± 0.0491 0.8359 ± 0.0283 -1.8650 ± 0.0746 0.6990 ± 0.0404
CPCP 0.7289 ± 0.0484 0.3729 ± 0.0151 0.1257 ± 0.0352 -2.0808 ± 0.3827 -2.2192 ± 0.0543 0.8707 ± 0.0487 -1.7002 ± 0.1060 0.7146 ± 0.0761
CPCP (Mix) 0.7298 ± 0.0538 0.3760 ± 0.0659 0.1316 ± 0.0329 -1.9282 ± 0.4194 -2.2567 ± 0.0470 0.8720 ± 0.0456 -1.7325 ± 0.0936 0.6744 ± 0.0550
CPCP (Clip) 0.7224 ± 0.0379 0.3673 ± 0.0273 0.1296 ± 0.0306 -2.2154 ± 0.4193 -2.2158 ± 0.0408 0.8603 ± 0.0346 -1.7748 ± 0.0854 0.6569 ± 0.0498
CPCP (Clip+Mix) 0.7384 ± 0.0683 0.3490 ± 0.0117 0.1258 ± 0.0369 -2.1888 ± 0.3348 -2.2360 ± 0.0359 0.8455 ± 0.0254 -1.7974 ± 0.0713 0.6189 ± 0.0465
Table 7: Comparison of MSCE ($K=50$). Mean Squared Coverage Error (MSCE, $\downarrow$)

### v3

@@ A.5 Detailed results
Table 5: Full comparison of MSCE ($K=10$) . MSCE ($\downarrow$) on all benchmarks with all baselines.
@@
Split 0.0031 ± 0.0008 0.0118 ± 0.0035 0.0033 ± 0.0006 0.0351 ± 0.0064 0.0039 ± 0.0010 0.0082 ± 0.0010 0.0125 ± 0.0031 0.0123 ± 0.0019
PLCP (G=20) 0.0021 ± 0.0011 0.0032 ± 0.0033 0.0008 ± 0.0009 0.0162 ± 0.0087 0.0026 ± 0.0013 0.0017 ± 0.0008 0.0027 ± 0.0038 0.0031 ± 0.0033
PLCP (G=50) 0.0016 ± 0.0007 0.0016 ± 0.0022 0.0008 ± 0.0010 0.0124 ± 0.0075 0.0015 ± 0.0015 0.0015 ± 0.0009 0.0009 ± 0.0006 0.0017 ± 0.0015
Gaussian-Scoring 0.0011 ± 0.0004 0.0006 ± 0.0003 0.0004 ± 0.0002 0.0129 ± 0.0095 0.0041 ± 0.0041 0.0026 ± 0.0011 0.0016 ± 0.0014 0.0057 ± 0.0026
CQR 0.0011 ± 0.0007 0.0010 ± 0.0004 0.0006 ± 0.0002 0.0120 ± 0.0064 0.0012 ± 0.0005 0.0009 ± 0.0005 0.0016 ± 0.0009 0.0061 ± 0.0007
CQR-ALD 0.0008 ± 0.0003 0.0007 ± 0.0007 0.0006 ± 0.0003 0.0055 ± 0.0059 0.0009 ± 0.0006 0.0015 ± 0.0010 0.0021 ± 0.0029 0.0045 ± 0.0009
RCP 0.0010 ± 0.0005 0.0013 ± 0.0006 0.0006 ± 0.0004 0.0029 ± 0.0020 0.0007 ± 0.0003 0.0013 ± 0.0011 0.0009 ± 0.0003 0.0030 ± 0.0005
RCP-ALD 0.0013 ± 0.0010 0.0010 ± 0.0006 0.0005 ± 0.0003 0.0080 ± 0.0102 0.0009 ± 0.0007 0.0011 ± 0.0005 0.0010 ± 0.0007 0.0025 ± 0.0008
RCP-MultiHead 0.0010 ± 0.0004 0.0015 ± 0.0006 0.0005 ± 0.0003 0.0027 ± 0.0013 0.0010 ± 0.0004 0.0011 ± 0.0008 0.0008 ± 0.0005 0.0031 ± 0.0008
CPCP 0.0009 ± 0.0004 0.0009 ± 0.0004 0.0004 ± 0.0002 0.0019 ± 0.0010 0.0003 ± 0.0002 0.0014 ± 0.0011 0.0009 ± 0.0006 0.0025 ± 0.0009
CPCP (Clip) 0.0010 ± 0.0004 0.0005 ± 0.0003 0.0005 ± 0.0003 0.0020 ± 0.0013 0.0003 ± 0.0002 0.0011 ± 0.0006 0.0007 ± 0.0004 0.0016 ± 0.0006
CPCP (Mix) 0.0009 ± 0.0005 0.0006 ± 0.0004 0.0004 ± 0.0002 0.0014 ± 0.0007 0.0004 ± 0.0001 0.0009 ± 0.0005 0.0006 ± 0.0003 0.0016 ± 0.0005
CPCP (Clip+Mix) 0.0008 ± 0.0002 0.0004 ± 0.0002 0.0004 ± 0.0002 0.0019 ± 0.0009 0.0003 ± 0.0001 0.0007 ± 0.0004 0.0004 ± 0.0002 0.0012 ± 0.0004
CPCP-$0.01$ 0.0010 ± 0.0004 0.0007 ± 0.0004 0.0005 ± 0.0002 0.0023 ± 0.0012 0.0002 ± 0.0001 0.0009 ± 0.0005 0.0010 ± 0.0006 0.0024 ± 0.0008
CPCP-$0.01$ (Clip+Mix) 0.0008 ± 0.0004 0.0004 ± 0.0003 0.0004 ± 0.0002 0.0012 ± 0.0007 0.0003 ± 0.0001 0.0006 ± 0.0004 0.0004 ± 0.0002 0.0013 ± 0.0005
CPCP-$0.05$ 0.0008 ± 0.0004 0.0008 ± 0.0005 0.0004 ± 0.0002 0.0017 ± 0.0009 0.0003 ± 0.0002 0.0009 ± 0.0004 0.0010 ± 0.0007 0.0026 ± 0.0013
CPCP-$0.05$ (Clip+Mix) 0.0007 ± 0.0003 0.0005 ± 0.0002 0.0004 ± 0.0002 0.0016 ± 0.0010 0.0003 ± 0.0002 0.0006 ± 0.0003 0.0004 ± 0.0003 0.0008 ± 0.0003
Table 6: Full comparison of MSCE ($K=30$). MSCE ($\downarrow$) on all benchmarks with all baselines.
@@
Split 0.0053 ± 0.0013 0.0138 ± 0.0021 0.0046 ± 0.0008 0.0392 ± 0.0070 0.0058 ± 0.0010 0.0089 ± 0.0012 0.0147 ± 0.0025 0.0186 ± 0.0016
PLCP (G=20) 0.0037 ± 0.0015 0.0045 ± 0.0041 0.0017 ± 0.0010 0.0200 ± 0.0078 0.0038 ± 0.0018 0.0040 ± 0.0012 0.0043 ± 0.0041 0.0062 ± 0.0055
PLCP (G=50) 0.0032 ± 0.0009 0.0025 ± 0.0036 0.0017 ± 0.0012 0.0155 ± 0.0082 0.0026 ± 0.0023 0.0040 ± 0.0016 0.0023 ± 0.0015 0.0036 ± 0.0024
Gaussian-Scoring 0.0023 ± 0.0006 0.0014 ± 0.0004 0.0010 ± 0.0003 0.0181 ± 0.0119 0.0056 ± 0.0052 0.0040 ± 0.0013 0.0026 ± 0.0020 0.0082 ± 0.0033
CQR 0.0025 ± 0.0012 0.0016 ± 0.0004 0.0011 ± 0.0003 0.0159 ± 0.0083 0.0018 ± 0.0007 0.0018 ± 0.0008 0.0034 ± 0.0014 0.0084 ± 0.0009
CQR-ALD 0.0019 ± 0.0005 0.0012 ± 0.0008 0.0011 ± 0.0003 0.0085 ± 0.0082 0.0014 ± 0.0009 0.0025 ± 0.0012 0.0036 ± 0.0030 0.0076 ± 0.0010
RCP 0.0019 ± 0.0006 0.0021 ± 0.0006 0.0012 ± 0.0004 0.0061 ± 0.0026 0.0011 ± 0.0004 0.0023 ± 0.0012 0.0016 ± 0.0003 0.0046 ± 0.0007
RCP-ALD 0.0023 ± 0.0010 0.0017 ± 0.0006 0.0012 ± 0.0005 0.0114 ± 0.0106 0.0014 ± 0.0009 0.0021 ± 0.0006 0.0023 ± 0.0024 0.0038 ± 0.0010
RCP-MultiHead 0.0020 ± 0.0004 0.0022 ± 0.0006 0.0011 ± 0.0004 0.0063 ± 0.0021 0.0015 ± 0.0004 0.0021 ± 0.0007 0.0016 ± 0.0007 0.0046 ± 0.0009
CPCP 0.0021 ± 0.0008 0.0020 ± 0.0008 0.0010 ± 0.0003 0.0054 ± 0.0019 0.0007 ± 0.0002 0.0026 ± 0.0018 0.0019 ± 0.0009 0.0043 ± 0.0013
CPCP (Clip) 0.0022 ± 0.0006 0.0012 ± 0.0005 0.0011 ± 0.0005 0.0056 ± 0.0019 0.0007 ± 0.0002 0.0022 ± 0.0008 0.0018 ± 0.0014 0.0035 ± 0.0007
CPCP (Mix) 0.0019 ± 0.0007 0.0015 ± 0.0008 0.0009 ± 0.0003 0.0046 ± 0.0014 0.0007 ± 0.0001 0.0021 ± 0.0013 0.0013 ± 0.0005 0.0029 ± 0.0008
CPCP (Clip+Mix) 0.0018 ± 0.0005 0.0010 ± 0.0002 0.0011 ± 0.0003 0.0054 ± 0.0019 0.0007 ± 0.0002 0.0016 ± 0.0005 0.0010 ± 0.0002 0.0026 ± 0.0008
CPCP-$0.01$ 0.0023 ± 0.0007 0.0020 ± 0.0008 0.0012 ± 0.0003 0.0062 ± 0.0018 0.0006 ± 0.0002 0.0023 ± 0.0009 0.0025 ± 0.0020 0.0042 ± 0.0011
CPCP-$0.01$ (Clip+Mix) 0.0018 ± 0.0005 0.0011 ± 0.0004 0.0009 ± 0.0002 0.0043 ± 0.0014 0.0007 ± 0.0002 0.0017 ± 0.0007 0.0010 ± 0.0004 0.0026 ± 0.0006
CPCP-$0.05$ 0.0018 ± 0.0004 0.0021 ± 0.0012 0.0011 ± 0.0003 0.0058 ± 0.0022 0.0007 ± 0.0002 0.0020 ± 0.0009 0.0023 ± 0.0012 0.0045 ± 0.0018
CPCP-$0.05$ (Clip+Mix) 0.0018 ± 0.0004 0.0010 ± 0.0002 0.0010 ± 0.0003 0.0049 ± 0.0016 0.0007 ± 0.0002 0.0015 ± 0.0003 0.0010 ± 0.0003 0.0020 ± 0.0004
Table 7: Full comparison of WSC. WSC ($\uparrow$) on all benchmarks with all baselines.
@@
Split 0.8133 ± 0.0277 0.6480 ± 0.0206 0.8192 ± 0.0159 0.5428 ± 0.0446 0.7435 ± 0.0156 0.7712 ± 0.0266 0.6797 ± 0.0234 0.7623 ± 0.0195
PLCP (G=20) 0.8559 ± 0.0341 0.8068 ± 0.0680 0.8690 ± 0.0229 0.6837 ± 0.0623 0.7891 ± 0.0514 0.8529 ± 0.0268 0.8123 ± 0.0645 0.8413 ± 0.0369
PLCP (G=50) 0.8663 ± 0.0203 0.8498 ± 0.0555 0.8738 ± 0.0162 0.7127 ± 0.0828 0.8303 ± 0.0608 0.8585 ± 0.0243 0.8450 ± 0.0208 0.8604 ± 0.0199
Gaussian-Scoring 0.8611 ± 0.0278 0.8613 ± 0.0170 0.8804 ± 0.0166 0.6928 ± 0.0911 0.7943 ± 0.0646 0.8434 ± 0.0232 0.8429 ± 0.0355 0.8259 ± 0.0347
CQR 0.8641 ± 0.0245 0.8563 ± 0.0159 0.8801 ± 0.0217 0.6997 ± 0.0837 0.8393 ± 0.0201 0.8704 ± 0.0235 0.8175 ± 0.0253 0.8149 ± 0.0188
CQR-ALD 0.8696 ± 0.0246 0.8735 ± 0.0259 0.8807 ± 0.0156 0.7836 ± 0.0745 0.8513 ± 0.0364 0.8622 ± 0.0246 0.8363 ± 0.0227 0.8222 ± 0.0193
RCP 0.8849 ± 0.0240 0.8448 ± 0.0237 0.8752 ± 0.0158 0.8002 ± 0.0457 0.8627 ± 0.0149 0.8638 ± 0.0260 0.8515 ± 0.0196 0.8516 ± 0.0198
RCP-ALD 0.8743 ± 0.0250 0.8558 ± 0.0178 0.8872 ± 0.0174 0.7596 ± 0.0803 0.8597 ± 0.0331 0.8779 ± 0.0181 0.8561 ± 0.0226 0.8582 ± 0.0173
RCP-MultiHead 0.8911 ± 0.0191 0.8361 ± 0.0172 0.8732 ± 0.0164 0.7985 ± 0.0286 0.8484 ± 0.0162 0.8745 ± 0.0247 0.8529 ± 0.0226 0.8519 ± 0.0150
CPCP 0.8820 ± 0.0132 0.8589 ± 0.0216 0.8844 ± 0.0125 0.8169 ± 0.0408 0.8864 ± 0.0119 0.8733 ± 0.0251 0.8548 ± 0.0172 0.8462 ± 0.0266
CPCP (Clip) 0.8895 ± 0.0213 0.8682 ± 0.0173 0.8805 ± 0.0175 0.8205 ± 0.0357 0.8860 ± 0.0188 0.8701 ± 0.0176 0.8706 ± 0.0191 0.8604 ± 0.0219
CPCP (Mix) 0.8812 ± 0.0182 0.8676 ± 0.0119 0.8816 ± 0.0141 0.8335 ± 0.0333 0.8873 ± 0.0139 0.8767 ± 0.0378 0.8736 ± 0.0137 0.8648 ± 0.0150
CPCP (Clip+Mix) 0.8882 ± 0.0250 0.8802 ± 0.0099 0.8912 ± 0.0139 0.8320 ± 0.0304 0.8912 ± 0.0126 0.8771 ± 0.0220 0.8759 ± 0.0158 0.8715 ± 0.0141
CPCP-$0.01$ 0.8830 ± 0.0141 0.8571 ± 0.0165 0.8794 ± 0.0126 0.8048 ± 0.0460 0.8852 ± 0.0094 0.8722 ± 0.0256 0.8581 ± 0.0229 0.8560 ± 0.0184
CPCP-$0.01$ (Clip+Mix) 0.8918 ± 0.0158 0.8783 ± 0.0200 0.8894 ± 0.0149 0.8390 ± 0.0357 0.8853 ± 0.0083 0.8801 ± 0.0199 0.8811 ± 0.0135 0.8700 ± 0.0211
CPCP-$0.05$ 0.8888 ± 0.0150 0.8617 ± 0.0202 0.8831 ± 0.0154 0.8252 ± 0.0399 0.8873 ± 0.0128 0.8688 ± 0.0227 0.8519 ± 0.0216 0.8435 ± 0.0254
CPCP-$0.05$ (Clip+Mix) 0.8915 ± 0.0162 0.8801 ± 0.0178 0.8841 ± 0.0123 0.8258 ± 0.0368 0.8865 ± 0.0129 0.8846 ± 0.0144 0.8809 ± 0.0145 0.8750 ± 0.0172
Table 8: Full comparison of $\ell_{1}$-ERT. $\ell_{1}$-ERT ($\downarrow$) on all benchmarks with all baselines.